# Revealing Fano Resonance in Dirac Materials ZrTe5 through Raman Scattering

### v2

@@ IV Theoretical calculations and analysis
In the above calculations, we assume perfect stoichiometry of an ideal crystal, In experiments, ZrTe5 samples exhibit some degree of electronic heterogeneity and nanostrip junctions  Kim et al. (2021), and excess Te incorporation into the sample during the annealing process can introduce n-type doping Chi et al. (2017). Therefore, we also consider how the temperature dependence of $D_{\textrm{e-h}}(\omega,T)$ can be affected by small electron doping. Generally, we find that electron doping shifts the peak of $D_{\textrm{e-h}}(\omega,T)$ toward higher temperature. At doping $0.06$ electron in the primitive cell with Fermi level shifted as shown in Fig. 4(a), $D_{\textrm{e-h}}(\omega,T)$ monotonically grows with temperature $T>100$K, as shown in the orange solid line in Fig. 4(c). This analysis implies the electronic heterogeneity in the sample can be an important factor for the experimentally observed Fano resonance enhancement in all the temperature region. We expect that newly developed terahertz nano-imaging methods Kim et al. (2023b, 2024, 2021), especially when applied at low temperatures and in strong magnetic fieldsKim et al. (2023a), will play a pivotal role in pushing forward the ongoing research. Nevertheless, the comprehensive analysis and explanation is much more involved, which we defer to future work,
@@
Figure 4: (a) Temperature-dependent Fermi Level without and with electron doping, shown in blue and orange curve. Doping electron is 0.06 e-. The temperature-dependent VBM and CBM at $\Gamma$ are showing in red and green curve. (b)Temperature-dependent joint electron–hole pair density of states $D_{e-h}(\omega,T)$ without (b) and with electron doping (c) shown in blue and orange curve, respectively.

### v3

@@ IV Theoretical calculations and analysis
In the above calculations, we assume perfect stoichiometry of an ideal crystal, In experiments, ZrTe5 samples exhibit some degree of electronic heterogeneity and nanostrip junctions  Kim et al. (2021), and excess Te incorporation into the sample during the annealing process can introduce n-type doping Chi et al. (2017). Therefore, we also consider how the temperature dependence of $D_{\textrm{e-h}}(\omega,T)$ can be affected by small electron doping. Our samples were synthesized using a similar Te-rich flux method as in Ref. Chi et al. (2017), where excess Te was shown to introduce n-type doping. Consistent with that study, our transport measurements revealed negative Hall and Seebeck coefficients at low temperatures, indicating the presence of n-type carriers. Generally, we find that electron doping shifts the peak of $D_{\textrm{e-h}}(\omega,T)$ toward higher temperature. At doping $0.06$ electron in the primitive cell with Fermi level shifted as shown in Fig. 4(a), $D_{\textrm{e-h}}(\omega,T)$ monotonically grows with temperature $T>100$ K, as shown in the red solid line in Fig. 4(c). This analysis implies the electronic heterogeneity in the sample can be an important factor for the experimentally observed Fano resonance enhancement in all the temperature region. We expect that newly developed terahertz nano-imaging methods Kim et al. (2023b, 2024, 2021), especially when applied at low temperatures and in strong magnetic fieldsKim et al. (2023a), will play a pivotal role in pushing forward the ongoing research.
The above analysis focuses on the key contribution of joint density of states $D_{\textrm{e-h}}(\omega,T)$ to the temperature-dependence of Fano resonance. However, the electron-phonon coupling strength $V_{\mathrm{e-ph}}$ in Eq. (4) may also depends on temperature and contribute to the variation of $1/q^{2}$ with temperature. Note that $V_{\mathrm{e-ph}}$ can often be estimated from the first derivative of the band energy position or band gap with respect to the phonon eigenmode displacement $\lambda$ Dicks et al. (2024); Yin et al. (2013). However, calculations for the B2g I mode reveal that the band gap exhibits a symmetric dependence on $\lambda$, with negligible linear terms near equilibrium ($\lambda=0$). This leads to a vanishing first-order derivative, indicating that a simple linear model cannot capture the coupling strength. Such symmetric behavior is reminiscent of infrared phonon modes in Ta2NiSe5, where the band gap renormalization shows a quadratic dependence on $\lambda$ as described by a two-band model Michael et al. (2024).
In contrast, the calculation results for ZrTe5 reveal an asymmetric energy shift between the valence and conduction band edges under B2g I phonon displacements, involving contributions from multiple bands (see Fig. 5 a-c). This complexity renders a two-band approximation insufficient, and a full multi-band analysis would be required for quantitatively extracting $V_{\text{e-ph}}$, which we defer to future work. Nevertheless, the overall magnitude of band shifts under finite phonon amplitude remains a meaningful indicator of the electron-phonon coupling strength, which we will use to analyze the trend of $V_{\text{e-ph}}$ with temperature. Fig. 5 presents the electronic band structures computed at selected temperatures (10 K, 135 K, and 300 K) under B2g \@slowromancapi@ mode displacements of $\lambda$ = 0.03 Å and 0.06 Å (Tez atom displacement). The resulting band shifts, shown in panels (a)–(c), highlight the impact of increasing phonon amplitude on the electronic structure. The black solid lines represent the unperturbed bands ($\lambda=0$), while the blue and red dashed lines correspond to distortions of $\lambda=\pm 0.03$ and ±$\pm 0.06$, where the band shift is found to be symmetric with respect to ±$\pm\lambda$. Panels (d) and (e) display the corresponding conduction band minimum (CBM, orange) and valence band maximum (VBM, green) energy shifts for $\lambda$ = 0.03 Å and 0.06 Å, respectively. The band shifts increase with temperature up to 135 K where the band gap closes, suggesting a temperature-enhancement for the electron-phonon coupling in this regime. Above 135 K, the band shift saturates, indicating a much weaker temperature-dependence of $V_{\text{e-ph}}$ at higher temperature. This analysis shows that $V_{\text{e-ph}}$ may contribute to the increase of $1/q^{2}$ at low temperature regime ($T<135K$), complementary to the contribution of $D_{\textrm{e-h}}(\omega,T)$ at higher temperatures ($T<135K$) when finite doping is considered.
@@
Figure 4: (a) Temperature-dependent Fermi Level without and with electron doping, shown in blue and red curve. Doping electron is 0.06 e-. The temperature-dependent VBM and CBM at $\Gamma$ are showing in green and orange curve. (b)Temperature-dependent joint electron–hole pair density of states $D_{e-h}(\omega,T)$ without (b) and with electron doping (c) shown in blue and red curve, respectively.
Figure 5: Temperature-dependent band structure shifts induced by B2g I phonon mode displacements. Panels (a)–(c) display the electronic band structures at 10 K, 135 K, and 300 K, respectively, for structures subjected to B2g \@slowromancapi@ phonon distortions. The black solid lines represent the unperturbed (undistorted) structures, while the blue and red dashed lines correspond to displacement amplitudes of $\lambda$ = 0.03 Å and 0.06 Å, respectively. The valence band maximum for $\lambda=0$ is shifted to zero for convenience of comparison across different temperatures. Panels (d) and (e) display the corresponding conduction band minimum (CBM, orange) and valence band maximum (VBM, green) energy shifts for larger amplitudes of $\lambda$ = 0.03 Å and 0.06 Å, respectively.